OpenStudy (anonymous):

What is the sum of one over six plus two over three in simplest form? one over six one over three two over three five over six

OpenStudy (calculusxy):

To do this find the least common denominator. In this case, it is 6. So the denominator of 2/3 needs to be changed to 4/6 because 2*2 is 4 and 3*2 is 6. 1/6 doesn't need to be changed because it already has 6 as the denominator. When you add it all up, you get 5/6. And this fraction can't be reduced.

OpenStudy (anonymous):

1/6+ 2/3= 1/6+ 4/6 =(1+4)/6=5/6